OpenStudy (anonymous):

Find the center and radius x^2 -18x +53 = 18y - y^2

OpenStudy (anonymous):

@pitamar do you know this?

OpenStudy (anonymous):

You're having here an equation of an ellipse. In order to solve it you first have to convert it to this form: $$ (ax-h)^2 + (by-v)^2 = r^2 $$

OpenStudy (anonymous):

we know that: \((a + b)^2 = a^2 + 2ab + b^2\) So we can tell: $$ x^2 - 18x + 53 = x^2 + 2 \cdot (-9)x + 53 = \\ = x^2 + 2 \cdot (-9)x + (-9)^2 - (-9)^2 + 53 = \\ =(x - 9)^2 - 81 + 53 = (x-9)^2 - 28 $$

OpenStudy (anonymous):

Can you do the same for the \(y\) on the right side?

OpenStudy (anonymous):

That whole equation just confused me

OpenStudy (anonymous):

I though the variables for a circle are h and k

OpenStudy (anonymous):

hehe, lol don't worry about that too much. If you're dealing with circles and not ellipse then you can rid off the \(a,b\) as well and you have: $$(x-h)^2 + (y-v)^2 = r^2$$

OpenStudy (anonymous):

in our case we have \((x-9)^2\) so \(h=9\)

OpenStudy (anonymous):

That looks better Lol

OpenStudy (anonymous):

ye I guess =p Ok, so can you complete the square of y on the right as well?

OpenStudy (anonymous):

Wait how did you get 9 for h

OpenStudy (anonymous):

on the equation we have \((x-h)^2\) and so far we have \((x-9)^2\) Don't worry about that yet, I just wanted to show you that we're getting the equation to this shape. Now you have to complete the square for y too and we can solve it =)

OpenStudy (anonymous):

Would y=9 too?

OpenStudy (anonymous):

I mean k

OpenStudy (anonymous):

ye seems so, but it's not enough, because in order to complete the square you have to add & subtract something, and that will remain in the equation afterwards. You can see that for doing it for the x I had to add \( (-9)^2 - (-9)^2 \) in order to form it into a square and then I was left with that \(-(-9)^2\)

OpenStudy (anonymous):

So would you need to get x and y on the left side of the equation

OpenStudy (anonymous):

You can move the y's to the left first if you want and then complete the square, that's fine

OpenStudy (anonymous):

You can also complete the square the way it is now on the right, whatever you prefer

OpenStudy (anonymous):

What is missing on the right in order to complete a square for y?

OpenStudy (anonymous):

Oh a number with a variable?

OpenStudy (anonymous):

without

OpenStudy (anonymous):

Ye, some number Let me write what we have and move the y's to the left it'll be easier: $$(x-9)^2 - 28 = 18y - y^2 \\ (x-9)^2 + y^2 - 18y = 28 $$

OpenStudy (anonymous):

We can add a number to both sides. what number is missing in order for us to be able to form a square?

OpenStudy (anonymous):

Oh I see what you mean now Lol. You need to subtract the 28 over and make it equal (x−9)2+y2−18y -28 = 0

OpenStudy (anonymous):

make it equal 0 forget the giberish

OpenStudy (anonymous):

Well the calculation is true but we don't want to get it equal to 0 We want to complete the square. so we get \((y + \text{something})^2\) instead of the \(y^2 - 18y\) we have now

OpenStudy (anonymous):

Just remember that if we had a square then it would be of the kind \((y + n)^2 = y^2 + 2ny + n^2\) We have to find what \(n\) is in our case in order to add \(n^2\) to both sides and then we'll be able to form the square \((y + n)^2\)

OpenStudy (anonymous):

want a hint?

OpenStudy (anonymous):

Yea that would help Lol I'm so confused

OpenStudy (anonymous):

Ok ok, We can see that we have \(y^2 - 18y\) we can form it as \(y^2 + 2 (-9)y\) which reveals us that the third term has to be \((-9)^2\) in order to complete the square to \((y+ (-9))^2 = (y-9)^2\) So we add \((-9)^2 = 81\) to both sides: $$ (x-9)^2 + y^2 -18y = 28 \\ (x-9)^2 + y^2 -18y + 81 = 28 + 81 \\ (x-9)^2 + (y-9)^2 = 109 $$

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

Oh yea now it does xD Thanks for the help bro. You saved me

OpenStudy (anonymous):

lol no problem ,glad I could help =)